# Say More with Less: Variable-Frame-Rate Speech Tokenization via Adaptive Clustering and Implicit Duration Coding

### v1

@@ Abstract
Existing speech tokenizers typically assign a fixed number of tokens per second, regardless of the varying information density or temporal fluctuations in the speech signal. This uniform token allocation mismatches the intrinsic structure of speech, where information is distributed unevenly over time. To address this, we propose VARSTok, a VAriable-frame-Rate Speech Tokenizer that adapts token allocation based on local feature similarity. VARSTok introduces two key innovations: (1) a temporal-aware density peak clustering algorithm that adaptively segments speech into variable-length units, and (2) a novel implicit duration coding scheme that embeds both content and temporal span into a single token index, eliminating the need for auxiliary duration predictors. Extensive experiments show that VARSTok significantly outperforms strong fixed-rate baselines. Notably, it achieves superior reconstruction naturalness while using up to 23% fewer tokens than a 40 Hz fixed-frame-rate baseline. VARSTok further yields lower word error rates and improved naturalness in zero-shot text-to-speech synthesis. To the best of our knowledge, this is the first work to demonstrate that a fully dynamic, variable-frame-rate acoustic speech tokenizer can be seamlessly integrated into downstream speech language models. Speech samples are available at https://zhengrachel.github.io/VARSTok.
@@ 1 Introduction
Speech tokenization has become central to modern speech modeling, powering advances in neural audio codecs (Zeghidour et al. 2021; Wu et al. 2023; Ai et al. 2024; Huang, Meng, and Ko 2024), generative speech synthesis (Zhang et al. 2024; Chen et al. 2025b), and multimodal large language models (LLMs) (Du et al. 2024b; Chen et al. 2025a). By converting continuous speech signals into discrete token sequences, speech tokenization bridges the gap between raw waveform and token-based language modeling architectures, thus enabling the application of LLMs to speech data.
Existing speech tokenization methods generally fall into three types: semantic tokenizers (Baevski et al. 2020; Hsu et al. 2021) that capture high-level linguistic content; acoustic tokenizers (Kumar et al. 2023; Xin et al. 2024; Liu et al. 2025) that preserve fine-grained signal fidelity; and hybrid approaches (Zhang et al. 2024; Ye et al. 2025a) that combine the merits of both. While recent acoustic tokenizers have shown impressive results, they typically operate at fixed frame rates (e.g., 75 Hz), uniformly allocating tokens across time, ignoring the underlying content or information density.
However, natural speech is inherently non-uniform over time (Keshishian, Norman-Haignere, and Mesgarani 2021). Segments with silence or stable vowels are acoustically redundant, while others with rapid articulatory transitions or expressive prosody carry dense information (Van Kuyk, Kleijn, and Hendriks 2017; Dieleman et al. 2021). Fixed-rate tokenizers fails to adapt to this variability, leading to inefficient token usage and poor alignment with temporal dynamics. This misalignment not only results in suboptimal compression but also hinders the ability of downsteam speech language models (LMs) to learn natural prosody and rhythm. Addressing this challenge requires a paradigm shift from fixed-rate to content-aware, dynamic tokenization.
To this end, we propose VARSTok, a fully dynamic, variable-frame-rate speech tokenizer that adaptively allocates tokens based on local feature similarity. It introduces a temporal-aware clustering algorithm to segment speech into variable-length units. To support downstream language modeling without auxiliary duration predictors, we design an implicit duration coding scheme that embeds both content and duration into a single token index. In contrast to prior work (Zhang et al. 2025), VARSTok enables fully dynamic token allocation without the need for hierarchical fusion or predefined temporal resolutions. Crucially, the resulting duration-aware tokens can be seamlessly used in autoregressive speech LMs without modification.
@@ 2.1 Speech Tokenization
Speech tokenization has gained increasing attention for its role in neural audio compression, generative speech modeling, and multimodal LLMs (Cui et al. 2024). Existing methods broadly fall into three categories.
Semantic tokenizers extract high-level linguistic units by applying clustering or vector quantization (VQ) (Van Den Oord, Vinyals et al. 2017) over features learned from large-scale self-supervised pretraining (Hsu et al. 2021; Baevski et al. 2020; Chen et al. 2022). While effective for speech understanding tasks (Yoon, Woo, and Kim 2024; Sharma 2022; Fang et al. 2025), they discard fine-grained acoustic details and do not support direct waveform reconstruction, limiting their utility in generative applications such as speech synthesis or editing.
Acoustic tokenizers focus on preserving waveform fidelity by mapping raw speech to discrete tokens using residual VQ (Zeghidour et al. 2021; Défossez et al. 2023; Kumar et al. 2023) Compared to semantic tokens, acoustic tokens retain richer speech details and can be directly used in generation pipelines without a separate token-to-waveform model. Recent approaches have focused on single-codebook designs (Xin et al. 2024; Zhai et al. 2025; Zeng et al. 2025) to improve token efficiency and compatibility with downstream speech LMs (Borsos et al. 2023a, b). WavTokenizer (Ji et al. 2025) achieves strong reconstruction quality at fixed frame rates of 40 Hz or 75 Hz using a single codebook. However, they still rely on fixed-rate token assignment, failing to account for the temporal variability of speech signals.
Hybrid tokenizers aim to combine the merits of both by distilling linguistic content into acoustic embeddings (Zhang et al. 2024; Ye et al. 2025a, b). While enabling high-quality synthesis with linguistic control, they often require multiple codebooks, hierarchical token streams, or auxiliary modules, increasing system complexity and reducing generality.
@@ 2.2 Adaptive Compression in Speech
Recent work has explored dynamic speech compression primarily in the context of semantic representation learning. For example, SD-HuBERT (Cho et al. 2024) uses sentence-level self-distillation to induce syllabic organization. Other approaches employ learnable temporal pooling (Dieleman et al. 2021), reinforcement learning (Cuervo et al. 2022), or syllable-aligned self-supervised representations (Baade, Peng, and Harwath 2025; Cho et al. 2025) for semantic boundary discovery. However, these methods are designed for semantic abstraction and discard fine-grained acoustic information, resulting in non-reconstructable representations that are inappropriate for waveform generation, transmission, or any task requiring high-fidelity reconstruction.
High-fidelity acoustic tokenizers predominantly, on the other hand, largely adopt a fixed-frame-rate paradigm, leaving dynamic alternatives underexplored. The most related work to ours is TFC (Zhang et al. 2025), which introduces variable temporal resolution into neural speech codec built upon multi-codebook residual VQ. TFC adaptively select among three predefined frame rates (75 Hz, 37.5 Hz, 18.75 Hz) based on entropy-based information density estimation, enabling flexible bitrate control and improved reconstruction under constrained token budgets. However, TFC relies on a fixed set of temporal granularities and constructs final representations by hierarchically fusing features across coarse-to-fine scales. Although the resulting frame rate varies, each token still adheres to one of the predefined resolutions, making the system pseudo-dynamic. Moreover, TFC ignores the modeling of token duration, limiting its utility in autoregressive modeling or alignment-sensitive tasks.
@@ 3.1 Overview
As illustrated in Figure 1, VARSTok comprises four major components: a speech encoder, a temporal-aware density peak clustering module, a VQ module, and a speech decoder. First, the speech encoder transforms raw waveforms into frame-level embeddings $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times H}$, where $T$ is the number of frames determined by the encoder’s base frame rate, and $H$ is the embedding dimension. Adjacent frames are then adaptively grouped into $N$ variable-length clusters $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{N}\}$ based on local feature similarity, using the temporal-aware density peak clustering module. Each cluster $\mathcal{C}_{n}$ is summarized into a mean-pooled embedding $\mathbf{z}_{n}\in\mathbb{R}^{H}$and quantized using a VQ module with a single codebook $\mathcal{E}=\{\mathbf{e}_{0},\dots,\mathbf{e}_{K-1}\}$ of size $K$, where $\mathbf{e}_{k}$ denotes the $k$-th codebook entry. The quantized embedding $\hat{\mathbf{z}}_{n}$ for each cluster is paired with its duration $d_{n}$, representing the number of original frames spaned by the cluster. This duration is further integrated into the token index by the proposed implicit duration coding scheme, allowing each token to jointly encode both content and temporal span, thereby preserving alignment without requiring auxiliary duration predictors. Before decoded by a speech decoder, the quantized embeddings $\hat{\mathbf{Z}}=[\hat{\mathbf{z}}_{1},\dots,\hat{\mathbf{z}}_{N}]$ are expanded according to their durations $\mathbf{d}=[d_{1},\dots,d_{N}]$ to form a frame-level sequence $\hat{\mathbf{Z}}^{R}=[\hat{\mathbf{z}}^{R}_{1},\dots,\hat{\mathbf{z}}^{R}_{T}]\in\mathbb{R}^{T\times H}$, restoring the temporal structure for high-fidelity waveform reconstruction.
@@ 3.2 Model Architecture
VARSTok adopts an encoder-VQ-decoder architecture following the design of WavTokenizer (Ji et al. 2025), which achieves advanced performance with a single codebook.
@@
The speech encoder begins with a 1D convolutional layer followed by four convolutional blocks, each containing a residual unit and a strided convolution for temporal downsampling while doubling the number of channels at each stage. The downsampled features are further processed by a two-layer bidirectional LSTM and a projection layer, producing frame-level embeddings $\mathbf{X}$. ELU activations are used throughout. The speech decoder adopts an improved architecture combining attention layers, ConvNextV2 (Woo et al. 2023) blocks, and an inverse Fourier transform-based upsampling module. In VARSTok, the decoder receives an expanded sequence of quantized cluster embeddings $\hat{\mathbf{Z}}^{R}$ obtained by repeating each quantized cluster embedding $\hat{\mathbf{z}}_{n}$ according to their durations $d_{n}$, and finally reconstructs the original waveform.
@@ 3.3 Temporal-Aware Density Peak Clustering
To adaptively segment the encoder output $\mathbf{X}$ into variable-length units, we propose a temporal-aware density peak clustering algorithm. Unlike prior clustering-based methods (Wu et al. 2025) designed for unordered inputs like image patches, speech features are inherently sequential and clustering must respect temporal continuity. Our proposed algorithm enforces strict time-order constraints to produce segments that are both semantically meaningful and temporally aligned. The full algorithm is detailed in Algorithm 1.
@@ 3.4 Implicit Duration Coding via Extended Index
A key challenge in variable-frame-rate speech tokenization is managing the duration of each token. Accurate duration modeling is critical for faithfully reconstructing the temporal structure of speech and ensuring proper alignment with other modalities in downstream speech LMs. A natural solution is to incorporate a separate FastSpeech-style duration predictor (Ren et al. 2019, 2020). However, this increases architectural complexity and breaks the simplicity of an end-to-end tokenizer. More importantly, we found emprically that jointly training a learned duration predictor within the tokenizer pipeline leads to severe optimization instability and poor convergence. To address this, we propose a simple yet effective implicit duration coding scheme that embeds both content identity and temporal span into a single token index. This design preserves temporal alignment without requiring auxiliary predictors, while maintaining the stability and modularity of the overall tokenizer architecture.
@@ 4.1 Experimental Setup
We trained VARSTok on the 585-hour LibriTTS corpus (Zen et al. 2019) at a 24 kHz sampling rate for fair comparison with prior work. Our experiments cover three distinct tasks for comprehensive evaluation:
@@
  • Semantic Representation Quality: To evaluate the semantic quality of the learned representations, we adopted the speech portion of the ARCH benchmark [^1] https://github.com/MorenoLaQuatra/ARCH. (La Quatra et al. 2024), including SLURP (Bastianelli et al. 2020) (intent classification), EMOVO (Costantini et al. 2014) and RAVDESS (Livingstone and Russo 2018) (emotion classification), and AudioMNIST (Becker et al. 2024) (spoken digit recognition). Additional details about the ARCH bechmark are provided in the Appendix E.
  • Downstream TTS Language Modeling: We evaluated VARSTok’s compatibility with downstream speech LMs by training an TTS model based on a decoder-only autoregressive architecture using LibriTTS dataset. Training details are provided in the Appendix F. Its zero-shot TTS performance was assessed on the test set from UniCATS (Du et al. 2024a).
@@
To further situate our model’s reconstruction performance within the broader field, we benchmarked it against several other representative tokenizers specifically on the speech reconstruction task. These include the single-codebook DAC (Kumar et al. 2023) with results quoted from the WavTokenizer paper, alongside BigCodec (Xin et al. 2024) and XCodec2.0 (Ye et al. 2025b) which we trained from scratch on 24 kHz LibriTTS corpus. We also included results from SQCodec as optimistic references. It is important to note that SQCodec and XCodec2.0 benefit from either being trained and evaluated on 16kHz speech using substantially larger datasets or leveraging additional semantic features during inference, making their results not directly comparable. For reproducibility, all detailed configurations for baseline models are provided in the Appendix D. We do not include earlier neural codecs such as EnCodec or SoundStream as baselines, as they rely on multi-codebook or hierarchical token streams that are not directly compatible with downstream speech language modeling. For experimental efficiency, we limited semantic evaluation and TTS comparisons to VARSTok and the 40 Hz variant of WavTokenizer.
@@
We employed a comprehensive set of standard metrics to evaluate VARSTok’s performance across three tasks. For speech reconstruction, we report UTMOS (Saeki et al. 2022) (a neural estimator of Mean-Opinion Score), PESQ (perceptual evaluation of speech quality), STOI (short-time objective intelligibility) and V/UV F1 Score (voiced/unvoiced frame classification accuracy) to assess the naturalness, quality, intelligibility, and prosodic consistency of the reconstructed speech, respectively. For VARSTok, the reported “Frame Rate” is the average token rate per second calculated across the entire test set. To ensure fair comparison, we calculate the bitrate for VARSTok based on the expanded token space as:
@@
For semantic evaluation, we use classification accuracy amd macro-averaged F1 score on the ARCH benchmark to quantify the discriminability of learned representations across the three tasks mentioned above. For TTS language modeling, we evaluated word error rate (WER) using the Whisper-large-v3 (Radford et al. 2023) model, speaker similarity via cosine similarity of WavLM-large [^3] https://huggingface.co/microsoft/wavlm-base-plus-sv. (Chen et al. 2022) embeddings, and UTMOS to assess perceptual naturalness of the synthesized speech. Subjective evaluations were also conducted. The setup is described in Appendix I. We report mean opinion scores (MOS) for naturalness and similarity MOS (SMOS) to assess perceived speaker similarity relative to the reference.

### v2

@@ Abstract
Existing speech tokenizers typically assign a fixed number of tokens per second, regardless of the varying information density or temporal fluctuations in the speech signal. This uniform token allocation mismatches the intrinsic structure of speech, where information is distributed unevenly over time. To address this, we propose VARSTok, a VAriable-frame-Rate Speech Tokenizer that adapts token allocation based on local feature similarity. VARSTok introduces two key innovations: (1) a temporal-aware density peak clustering algorithm that adaptively segments speech into variable-length units, and (2) a novel implicit duration coding scheme that embeds both content and temporal span into a single token index, eliminating the need for auxiliary duration predictors. Extensive experiments show that VARSTok significantly outperforms strong fixed-rate baselines. Notably, it achieves superior reconstruction naturalness while using up to 23% fewer tokens than a 40 Hz fixed-frame-rate baseline. VARSTok further yields lower word error rates and improved naturalness in zero-shot text-to-speech synthesis. To the best of our knowledge, this is the first work to demonstrate that a fully dynamic, variable-frame-rate acoustic speech tokenizer can be seamlessly integrated into downstream speech language models.
Project page with samples and codehttps://zhengrachel.github.io/VARSTok
Extended version with appendixhttps://arxiv.org/abs/2509.04685
@@ 1 Introduction
Speech tokenization has become central to modern speech modeling, powering advances in neural audio codecs (zeghidour2021soundstream; wu2023audiodec; ai2024apcodec; huang2024repcodec), generative speech synthesis (zhang2024speechtokenizer; chen2025neural), and multimodal large language models (LLMs) (du2024cosyvoice; chen2025minmo). By converting continuous speech signals into discrete token sequences, speech tokenization bridges the gap between raw waveform and token-based language modeling architectures, thus enabling the application of LLMs to speech data.
Existing speech tokenization methods generally fall into three types: semantic tokenizers (baevski2020wav2vec; hsu2021hubert) that capture high-level linguistic content; acoustic tokenizers (kumar2023high; xin2024bigcodec; liu2025analyzing) that preserve fine-grained signal fidelity; and hybrid approaches (zhang2024speechtokenizer; ye2025codec) that combine the merits of both. While recent acoustic tokenizers have shown impressive results, they typically operate at fixed frame rates (e.g., 75 Hz), uniformly allocating tokens across time, ignoring the underlying content or information density.
However, natural speech is inherently non-uniform over time (keshishian2021understanding). Segments with silence or stable vowels are acoustically redundant, while others with rapid articulatory transitions or expressive prosody carry dense information (van2017information; dieleman2021variable). Fixed-rate tokenizers fails to adapt to this variability, leading to inefficient token usage and poor alignment with temporal dynamics. This misalignment not only results in suboptimal compression but also hinders the ability of downsteam speech language models (LMs) to learn natural prosody and rhythm. Addressing this challenge requires a paradigm shift from fixed-rate to content-aware, dynamic tokenization.
To this end, we propose VARSTok, a fully dynamic, variable-frame-rate speech tokenizer that adaptively allocates tokens based on local feature similarity. It introduces a temporal-aware clustering algorithm to segment speech into variable-length units. To support downstream language modeling without auxiliary duration predictors, we design an implicit duration coding scheme that embeds both content and duration into a single token index. In contrast to prior work (zhang2025unlocking), VARSTok enables fully dynamic token allocation without the need for hierarchical fusion or predefined temporal resolutions. Crucially, the resulting duration-aware tokens can be seamlessly used in autoregressive speech LMs without modification.
@@ 2.1 Speech Tokenization
Speech tokenization has gained increasing attention for its role in neural audio compression, generative speech modeling, and multimodal LLMs (cui2024recent). Existing methods broadly fall into three categories.
Semantic tokenizers extract high-level linguistic units by applying clustering or vector quantization (VQ) (van2017neural) over features learned from large-scale self-supervised pretraining (hsu2021hubert; baevski2020wav2vec; chen2022wavlm). While effective for speech understanding tasks (yoon2024hubert; sharma2022multi; fang2024llama), they discard fine-grained acoustic details and do not support direct waveform reconstruction, limiting their utility in generative applications such as speech synthesis or editing.
Acoustic tokenizers focus on preserving waveform fidelity by mapping raw speech to discrete tokens using residual VQ (zeghidour2021soundstream; defossezhigh; kumar2023high) Compared to semantic tokens, acoustic tokens retain richer speech details and can be directly used in generation pipelines without a separate token-to-waveform model. Recent approaches have focused on single-codebook designs (xin2024bigcodec; zhai2025one; zengscaling) to improve token efficiency and compatibility with downstream speech LMs (borsos2023audiolm; borsos2023soundstorm). WavTokenizer (ji2024wavtokenizer) achieves strong reconstruction quality at fixed frame rates of 40 Hz or 75 Hz using a single codebook. However, they still rely on fixed-rate token assignment, failing to account for the temporal variability of speech signals.
Hybrid tokenizers aim to combine the merits of both by distilling linguistic content into acoustic embeddings (zhang2024speechtokenizer; ye2025codec; ye2025llasa). While enabling high-quality synthesis with linguistic control, they often require multiple codebooks, hierarchical token streams, or auxiliary modules, increasing system complexity and reducing generality.
@@ 2.2 Adaptive Compression in Speech
Recent work has explored dynamic speech compression primarily in the context of semantic representation learning. For example, SD-HuBERT (cho2024sd) uses sentence-level self-distillation to induce syllabic organization. Other approaches employ learnable temporal pooling (dieleman2021variable), reinforcement learning (cuervo2022variable), or syllable-aligned self-supervised representations (baadesyllablelm; chosylber) for semantic boundary discovery. However, these methods are designed for semantic abstraction and discard fine-grained acoustic information, resulting in non-reconstructable representations that are inappropriate for waveform generation, transmission, or any task requiring high-fidelity reconstruction.
High-fidelity acoustic tokenizers predominantly, on the other hand, largely adopt a fixed-frame-rate paradigm, leaving dynamic alternatives underexplored. The most related work to ours is TFC (zhang2025unlocking), which introduces variable temporal resolution into neural speech codec built upon multi-codebook residual VQ. TFC adaptively select among three predefined frame rates (75 Hz, 37.5 Hz, 18.75 Hz) based on entropy-based information density estimation, enabling flexible bitrate control and improved reconstruction under constrained token budgets. However, TFC relies on a fixed set of temporal granularities and constructs final representations by hierarchically fusing features across coarse-to-fine scales. Although the resulting frame rate varies, each token still adheres to one of the predefined resolutions, making the system pseudo-dynamic. Moreover, TFC ignores the modeling of token duration, limiting its utility in autoregressive modeling or alignment-sensitive tasks.
@@ 3.1 Overview
As illustrated in Figure 1, VARSTok comprises four major components: a speech encoder, a temporal-aware density peak clustering module, a VQ module, and a speech decoder. First, the speech encoder transforms raw waveforms into frame-level embeddings $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times H}$, where $T$ is the number of frames determined by the encoder’s base frame rate, and $H$ is the embedding dimension. Adjacent frames are then adaptively grouped into NN variable-length clusters $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{N}\}$ based on local feature similarity, using the temporal-aware density peak clustering module. Each cluster $\mathcal{C}_{n}$ is summarized into a mean-pooled embedding $\mathbf{z}_{n}\in\mathbb{R}^{H}$and quantized using a VQ module with a single codebook $\mathcal{E}=\{\mathbf{e}_{0},\dots,\mathbf{e}_{K-1}\}$ of size $K$, where $\mathbf{e}_{k}$ denotes the $k$-th codebook entry. The quantized embedding $\hat{\mathbf{z}}_{n}$ for each cluster is paired with its duration $d_{n}$, representing the number of original frames spaned by the cluster. This duration is further integrated into the token index by the proposed implicit duration coding scheme, allowing each token to jointly encode both content and temporal span, thereby preserving alignment without requiring auxiliary duration predictors. Before decoded by a speech decoder, the quantized embeddings $\hat{\mathbf{Z}}=[\hat{\mathbf{z}}_{1},\dots,\hat{\mathbf{z}}_{N}]$ are expanded according to their durations $\mathbf{d}=[d_{1},\dots,d_{N}]$ to form a frame-level sequence $\hat{\mathbf{Z}}^{R}=[\hat{\mathbf{z}}^{R}_{1},\dots,\hat{\mathbf{z}}^{R}_{T}]\in\mathbb{R}^{T\times H}$, restoring the temporal structure for high-fidelity waveform reconstruction.
@@ 3.2 Model Architecture
VARSTok adopts an encoder-VQ-decoder architecture following the design of WavTokenizer (ji2024wavtokenizer), which achieves advanced performance with a single codebook.
@@
The speech encoder begins with a 1D convolutional layer followed by four convolutional blocks, each containing a residual unit and a strided convolution for temporal downsampling while doubling the number of channels at each stage. The downsampled features are further processed by a two-layer bidirectional LSTM and a projection layer, producing frame-level embeddings $\mathbf{X}$. ELU activations are used throughout. The speech decoder adopts an improved architecture combining attention layers, ConvNextV2 (woo2023convnext) blocks, and an inverse Fourier transform-based upsampling module. In VARSTok, the decoder receives an expanded sequence of quantized cluster embeddings $\hat{\mathbf{Z}}^{R}$ obtained by repeating each quantized cluster embedding $\hat{\mathbf{z}}_{n}$ according to their durations $d_{n}$, and finally reconstructs the original waveform.
@@ 3.3 Temporal-Aware Density Peak Clustering
To adaptively segment the encoder output $\mathbf{X}$ into variable-length units, we propose a temporal-aware density peak clustering algorithm. Unlike prior clustering-based methods (wu2024towards) designed for unordered inputs like image patches, speech features are inherently sequential and clustering must respect temporal continuity. Our proposed algorithm enforces strict time-order constraints to produce segments that are both semantically meaningful and temporally aligned. The full algorithm is detailed in Algorithm 1.
@@ 3.4 Implicit Duration Coding via Extended Index
A key challenge in variable-frame-rate speech tokenization is managing the duration of each token. Accurate duration modeling is critical for faithfully reconstructing the temporal structure of speech and ensuring proper alignment with other modalities in downstream speech LMs. A natural solution is to incorporate a separate FastSpeech-style duration predictor (ren2019fastspeech; renfastspeech). However, this increases architectural complexity and breaks the simplicity of an end-to-end tokenizer. More importantly, we found emprically that jointly training a learned duration predictor within the tokenizer pipeline leads to severe optimization instability and poor convergence. To address this, we propose a simple yet effective implicit duration coding scheme that embeds both content identity and temporal span into a single token index. This design preserves temporal alignment without requiring auxiliary predictors, while maintaining the stability and modularity of the overall tokenizer architecture.
@@ 4.1 Experimental Setup
We trained VARSTok on the 585-hour LibriTTS corpus (zen2019libritts) at a 24 kHz sampling rate for fair comparison with prior work. Our experiments cover three distinct tasks for comprehensive evaluation:
@@
  • Semantic Representation Quality: To evaluate the semantic quality of the learned representations, we adopted the speech portion of the ARCH benchmark [^1] https://github.com/MorenoLaQuatra/ARCH. (la2024benchmarking), including SLURP (bastianelli2020slurp) (intent classification), EMOVO (costantini2014emovo) and RAVDESS (livingstone2018ryerson) (emotion classification), and AudioMNIST (becker2024audiomnist) (spoken digit recognition). Additional details about the ARCH bechmark are provided in the Appendix E.
  • Downstream TTS Language Modeling: We evaluated VARSTok’s compatibility with downstream speech LMs by training an TTS model based on a decoder-only autoregressive architecture using LibriTTS dataset. Training details are provided in the Appendix F. Its zero-shot TTS performance was assessed on the test set from UniCATS (du2024unicats).
@@
To further situate our model’s reconstruction performance within the broader field, we benchmarked it against several other representative tokenizers specifically on the speech reconstruction task. These include the single-codebook DAC (kumar2023high) with results quoted from the WavTokenizer paper, alongside BigCodec (xin2024bigcodec) and XCodec2.0 (ye2025llasa) which we trained from scratch on 24 kHz LibriTTS corpus. We also included results from SQCodec as optimistic references. It is important to note that SQCodec and XCodec2.0 benefit from either being trained and evaluated on 16kHz speech using substantially larger datasets or leveraging additional semantic features during inference, making their results not directly comparable. For reproducibility, all detailed configurations for baseline models are provided in the Appendix D. We do not include earlier neural codecs such as EnCodec or SoundStream as baselines, as they rely on multi-codebook or hierarchical token streams that are not directly compatible with downstream speech language modeling. For experimental efficiency, we limited semantic evaluation and TTS comparisons to VARSTok and the 40 Hz variant of WavTokenizer.
@@
We employed a comprehensive set of standard metrics to evaluate VARSTok’s performance across three tasks. For speech reconstruction, we report UTMOS (saeki2022utmos) (a neural estimator of Mean-Opinion Score), PESQ (perceptual evaluation of speech quality), STOI (short-time objective intelligibility) and V/UV F1 Score (voiced/unvoiced frame classification accuracy) to assess the naturalness, quality, intelligibility, and prosodic consistency of the reconstructed speech, respectively. For VARSTok, the reported “Frame Rate” is the average token rate per second calculated across the entire test set. To ensure fair comparison, we calculate the bitrate for VARSTok based on the expanded token space as:
@@
For semantic evaluation, we use classification accuracy amd macro-averaged F1 score on the ARCH benchmark to quantify the discriminability of learned representations across the three tasks mentioned above. For TTS language modeling, we evaluated word error rate (WER) using the Whisper-large-v3 (radford2023robust) model, speaker similarity via cosine similarity of WavLM [^3] https://huggingface.co/microsoft/wavlm-base-plus-sv. (chen2022wavlm) embeddings, and UTMOS to assess perceptual naturalness of the synthesized speech. Subjective evaluations were also conducted. The setup is described in Appendix I. We report mean opinion scores (MOS) for naturalness and similarity MOS (SMOS) to assess perceived speaker similarity relative to the reference.